# Designing Multi-layered Runtime Guardrails for Foundation Model Based Agents: Swiss Cheese Model for AI Safety by Design

### v3

@@ IV-A9 Customizability
The traceability attribute of guardrails tracks and records the origins, processes, and decision paths, such as input and output of FMs, external tools, etc. [27]. It involves maintaining detailed logs and records that can be audited to understand how decisions are made. For example, in a customer service chatbot, traceability ensures that every recommendation can be traced back to the data sources and algorithms used. This provides a clear audit trail for transparency and accountability. Traceability also aids in identifying the root causes of issues to enable timely and accurate troubleshooting and improvement [26], and helps in maintaining user trust and meeting regulatory requirements [16, 10]. Comprehensive documentation of data sources and model modifications also better support effective auditing and compliance checking [12].
@@ IV-B1 Actions
TABLE V: A Mapping of Agent Targets to Guardrail Actions
@@ IV-B2 Targets
Guardrail actions can be applied to various targets across multi-layers, including both pipelines and artifacts. Some guardrails are applied the the entire pipeline (including prompts, intermediate results, and final results), while others are focused on specific artifacts (covering goals, context, reasoning, plans, memory, tools, knowledge bases, other agents, FMs). Table V provides an overview of agent targets and corresponding guardrail actions.
@@
  • Tools: Overseeing the proper use of tools by agents, including implementing access controls, restricting tool capabilities, and detect potential vulnerabilities [49, 36].
@@ IV-B3 Rules
IV-B4 Sources
The source of guardrails in FM-based agents ranges from industry regulations and standards to individual preferences. Industry-level regulations and standards provide the broader regulatory framework within which FM-based agents must operate. Guardrails designed to comply with these regulations guarantee that the system adheres to industry best practices and legal requirements [16]. They facilitate simpler auditing and certification processes, ensuring the agent remains compliant with evolving regulatory landscapes.
@@ IV-B5 Modality
Multimodal guardrails address the combined risks of handling multiple data types. They synchronize protections across different data types, ensuring comprehensive security and compliance [55]. For example, a system that generates text based on image inputs must ensure accurate and ethical representation of the image content. This requires advanced cross-modal analysis and validation techniques to ensure the system operates reliably and ethically across all data types it handles [53].
@@ V Reference Architecture for Designing Multi-layered Runtime Guardrails of Agents
Figure 4 shows the proposed reference architecture for multi-layered runtime guardrails of FM-based agents, which consists of four key parts: (i) external environment, (ii) agent components, (iii) built-in multi-layered runtime guardrails, and (iv) AgentOps infrastructure.
@@ V-A External Environment:
The external environment refers to all entities interacting with the agent, including users, other agents, external tools, and knowledge bases. Users provide goals and contextual inputs that shape the agent’s objectives. To achieve user goals, the agent may utilize context detected in the external environment and interact with other agents, specialized tools, and extensive knowledge bases to perform complex tasks.
@@ V-B Agent Components
  • Context Engine: The context engine processes multimodal context data from the external environment to enrich the user prompt, helping FMs better understand user goals. A prompt may contain elements such as goals and context. Instead of waiting for users’ instructions, the agent can also proactively make suggestions based on the context it detects, such as screen recordings, mouse clicks, eye tracking data, gestures, and document annotations [3].
@@
  • Memory: The memory component in this architecture stores relevant information from prior interactions, plans, and results. This accumulated knowledge supports continuous learning, enabling the agent to refine its strategies and improve capabilities and skills over time, thereby improving accuracy and minimizing repeated errors.
@@ V-C Multi-layered Runtime Guardrails
Building on the Swiss Cheese Model, we design multi-layered runtime guardrails for FM-based agents, structured around the dimensions of quality attributes, pipelines, and artifacts specified in the taxonomy. In this architecture, each ‘cheese slice’ represents a protective layer within the agent system, addressing quality attributes, pipeline stages, and/or specific artifacts, such as a layer about privacy guardrails for prompts or security guardrails for tools. While each layer contains holes (i.e., potential gaps or weaknesses), where risks might slip through, the holes are positioned differently across layers. Gaps in one layer are often covered by another; thus, even if one layer fails, another can catch and mitigate the issue.
From the perspective of quality attributes (discussed in section IV-A), guardrails can be designed to ensure accuracy, efficiency, privacy, security, safety, fairness, compliance. From the pipelines perspective, guardrails can be applied at multiple stages: the user prompts, intermediate results during workflow executions, and final results generated by the agent.
@@
  • Guardrails for workflows: Handle the exceptions that arise during the workflow executions by implementing mechanisms like force-failing a step or retrying a tool call [80]
@@
  • Guardrails for FMs: Enforce boundaries on the FM’s non-deterministic outputs, applying modifications or flags as needed.
@@ VI Threats to Validity
Moreover, our definitions and categorizations may not capture all relevant aspects of guardrails in FM-based agents. To mitigate this threat, we validated the taxonomy through extensive literature review and expert feedback. However, this introduces a risk of producing biased results that address only expert needs, as the people involved in the feedback process have extensive experience in the AI and software engineering domains. Our review protocols helped us to reduce such bias.
We prepared a guardrails taxonomy and conducted a comparative analysis of its components to help the reader better understand their design and evaluation. We critically examined the strength and consistency of relationships in the selected studies to develop a reliable taxonomy and reference architecture for designing built-in multi-layered runtime guardrails. Finally, we draw conclusions. Nonetheless, the generalizability of guardrails to different contexts and types in FM-based agents remains a potential limitation. Specific adaptations might be necessary for certain systems, such as those used in healthcare or financial organizations.
@@ VII Conclusion and Future Work
To advance the understanding of runtime guardrail design in FM-based agents, this paper presents a comprehensive taxonomy of guardrail design based on the results of an SLR. Our taxonomy categorizes guardrails based on their essential quality attributes and key design dimensions, including guardrail actions and targets, employed rules, guardrail sources, modality, and underlying models. Building on this taxonomy, we propose a novel Swiss Cheese Model for AI safety - a reference architecture for designing built-in, multi-layered guardrails in FM-based agents, which includes three dimensions: quality attributes, pipelines, and artifacts. In the future, we plan to develop guardrail services for a scientific agent platform, implementing the proposed reference architecture and integrating various design options outlined in the taxonomy.
Appendix A: List of Selected Studies
  • SS [1] M. Pawagi, V. Kumar, Guardrails: Automated suggestions for clarifying ambiguous purpose statements, in: Proceedings of the 16th Annual ACM India Compute Conference, COMPUTE ’23, Association for Computing Machinery, New York, NY, USA, 2023, p. 55–60. doi:10.1145/3627217.3627234.
  • SS [2] A. Khorramrouz, S. Dutta, A. Dutta, A. R. KhudaBukhsh, Down the toxicity rabbit hole: Investigating palm 2 guardrails, arXiv preprint arXiv:2309.06415 (2023). doi:https://doi.org/10.48550/arXiv.2309.06415.
  • SS [3] Y. Dong, R. Mu, G. Jin, Y. Qi, J. Hu, X. Zhao, J. Meng, W. Ruan, X. Huang, Building guardrails for large language models, arXiv preprint arXiv:2402.01822 (2024). doi:https://doi.org/10.48550/arXiv.2402.01822.
  • SS [4] N. Mangaokar, A. Hooda, J. Choi, S. Chandrashekaran, K. Fawaz, S. Jha, A. Prakash, Prp: Propagating universal perturbations to attack large language model guard-rails, arXiv preprint arXiv:2402.15911 (2024). doi:https://doi.org/10.48550/arXiv.2402.15911.
  • SS [5] M. Anderljung, J. Barnhart, J. Leung, A. Korinek, C. O’Keefe, J. Whittlestone, S. Avin, M. Brundage, J. Bullock, D. Cass-Beggs, et al., Frontier ai regulation: Managing emerging risks to public safety, arXiv preprint arXiv:2307.03718 (2023). doi:https://doi.org/10.48550/arXiv.2307.03718.
  • SS [6] M. Liffiton, B. E. Sheese, J. Savelka, P. Denny, Codehelp: Using large language models with guardrails for scalable support in programming classes, in: Proceedings of the 23rd Koli Calling International Conference on Computing Education Research, Koli Calling ’23, Association for Computing Machinery, New York, NY, USA, 2024, pp. 1–11. doi:10.1145/3631802.3631830.
  • SS [7] Z. Zhang, Y. Lu, J. Ma, D. Zhang, R. Li, P. Ke, H. Sun, L. Sha, Z. Sui, H. Wang, et al., Shieldlm: Empowering llms as aligned, customizable and explainable safety detectors, arXiv preprint arXiv:2402.16444 (2024). doi:https://doi.org/10.48550/arXiv.2402.16444.
  • SS [8] T. Rebedea, R. Dinu, M. N. Sreedhar, C. Parisien, J. Cohen, NeMo guardrails: A toolkit for controllable and safe LLM applications with programmable rails, in: Y. Feng, E. Lefever (Eds.), Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Association for Computational Linguistics, Singapore, 2023, pp. 431–445. doi:10.18653/v1/2023.emnlp-demo.40.
  • SS [9] Y. Wang, L. Singh, Adding guardrails to advanced chatbots, arXiv preprint arXiv:2306.07500 (2023). doi:https://doi.org/10.48550/arXiv.2306.07500.
  • SS [10] M. Shanahan, Talking about large language models, Commun. ACM 67 (2) (2024) 68–79. doi:10.1145/3624724.
  • SS [11] W. Du, Q. Li, J. Zhou, X. Ding, X. Wang, Z. Zhou, J. Liu, Finguard: A multimodal aigc guardrail in financial scenarios, in: Proceedings of the 5th ACM International Conference on Multimedia in Asia, MMAsia ’23, Association for Computing Machinery, New York, NY, USA, 2024, pp. 1–3. doi:10.1145/3595916.3626351.
  • SS [12] A. Wei, N. Haghtalab, J. Steinhardt, Jailbroken: How does llm safety training fail?, Advances in Neural Information Processing Systems 36 (2024). doi:10.48550/arXiv.2307.02483.
  • SS [13] J. Zhao, K. Chen, X. Yuan, Y. Qi, W. Zhang, N. Yu, Silent guardian: Protecting text from malicious exploitation by large language models, arXiv preprint arXiv:2312.09669 (2023). doi:https://doi.org/10.48550/arXiv.2312.09669.
  • SS [14] X. Qi, Y. Zeng, T. Xie, P.-Y. Chen, R. Jia, P. Mittal, P. Henderson, Fine-tuning aligned language models compromises safety, even when users do not intend to!, arXiv preprint arXiv:2310.03693 (2023). doi:https://doi.org/10.48550/arXiv.2310.03693.
  • SS [15] J. Mökander, J. Schuett, H. R. Kirk, L. Floridi, Auditing large language models: a three-layered approach, AI and Ethics (2023) 1–31doi:https://doi.org/10.1007/s43681-023-00289-2.
  • SS [16] S. Banerjee, S. Layek, R. Hazra, A. Mukherjee, How (un) ethical are instruction-centric responses of llms? unveiling the vulnerabilities of safety guardrails to harmful queries, arXiv preprint arXiv:2402.15302 (2024). doi:https://doi.org/10.48550/arXiv.2402.15302.
  • SS [17] S. Ee—Researcher, Z. W. Director, Adapting cybersecurity frameworks to manage frontier AI risks, Institute for AI Policy and Strategy (IAPS) (2023).
    URL https://static1.squarespace.com/static/64edf8e7f2b10d716b5ba0e1/t/6528c5c7f912f74fbd03fc34/1697170896984/Adapting+cybersecurity+frameworks+to+manage+frontier+AI+risks.pdf
  • SS [18] P. Rai, S. Sood, V. K. Madisetti, A. Bahga, Guardian: A multi-tiered defense architecture for thwarting prompt injection attacks on llms, Journal of Software Engineering and Applications 17 (1) (2024) 43–68. doi:https://doi.org/10.4236/jsea.2024.171003.
  • SS [19] A. Kumar, S. Singh, S. V. Murty, S. Ragupathy, The ethics of interaction: Mitigating security threats in llms, arXiv preprint arXiv:2401.12273 (2024). doi:https://doi.org/10.48550/arXiv.2401.12273.
  • SS [20] X. Shen, Z. Chen, M. Backes, Y. Shen, Y. Zhang, ”do anything now”: Characterizing and evaluating in-the-wild jailbreak prompts on large language models, arXiv preprint arXiv:2308.03825 (2023). doi:https://doi.org/10.48550/arXiv.2308.03825.
  • SS [21] A. Kumar, C. Agarwal, S. Srinivas, S. Feizi, H. Lakkaraju, Certifying llm safety against adversarial prompting, arXiv preprint arXiv:2309.02705 (2023). doi:https://doi.org/10.48550/arXiv.2309.02705.
  • SS [22] Y. Zeng, H. Lin, J. Zhang, D. Yang, R. Jia, W. Shi, How johnny can persuade llms to jailbreak them: Rethinking persuasion to challenge ai safety by humanizing llms, arXiv preprint arXiv:2401.06373 (2024). doi:https://doi.org/10.48550/arXiv.2401.06373.
  • SS [23] B. Wei, K. Huang, Y. Huang, T. Xie, X. Qi, M. Xia, P. Mittal, M. Wang, P. Henderson, Assessing the brittleness of safety alignment via pruning and low-rank modifications, arXiv preprint arXiv:2402.05162 (2024). doi:https://doi.org/10.48550/arXiv.2402.05162.
  • SS [24] S. Goyal, M. Hira, S. Mishra, S. Goyal, A. Goel, N. Dadu, D. Kirushikesh, S. Mehta, N. Madaan, Llmguard: Guarding against unsafe llm behavior, in: Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38(21), 2024, pp. 23790–23792. doi:https://doi.org/10.1609/aaai.v38i21.30566.
  • SS [25] R. R. Llaca, V. Leskoschek, V. C. Paiva, C. Lupău, P. Lippmann, J. Yang, Student-teacher prompting for red teaming to improve guardrails, in: Proceedings of the ART of Safety: Workshop on Adversarial testing and Red-Teaming for generative AI, 2023, pp. 11–23. doi:http://dx.doi.org/10.18653/v1/2023.artofsafety-1.2.
  • SS [26] Z. Wang, F. Yang, L. Wang, P. Zhao, H. Wang, L. Chen, Q. Lin, K.-F. Wong, SELF-GUARD: Empower the LLM to safeguard itself, in: K. Duh, H. Gomez, S. Bethard (Eds.), Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), Association for Computational Linguistics, Mexico City, Mexico, 2024, pp. 1648–1668. URL https://aclanthology.org/2024.naacl-long.92
  • SS [27] B. Wang, W. Chen, H. Pei, C. Xie, M. Kang, C. Zhang, C. Xu, Z. Xiong, R. Dutta, R. Schaeffer, et al., Decodingtrust: A comprehensive assessment of trustworthiness in gpt models, in: Advances in Neural Information Processing Systems, 2023, pp. 1–110. doi:https://doi.org/10.48550/arXiv.2306.11698.
  • SS [28] R. Bommasani, D. A. Hudson, E. Adeli, R. B. Altman, S. Arora, S. von Arx, M. S. Bernstein, J. Bohg, A. Bosselut, E. Brunskill, E. Brynjolfsson, S. Buch, D. Card, R. Castellon, N. S. Chatterji, A. S. Chen, K. Creel, J. Q. Davis, D. Demszky, C. Donahue, M. Doumbouya, E. Durmus, S. Ermon, J. Etchemendy, K. Ethayarajh, L. Fei-Fei, C. Finn, T. Gale, L. Gillespie, K. Goel, N. D. Goodman, S. Grossman, N. Guha, T. Hashimoto, P. Henderson, J. Hewitt, D. E. Ho, J. Hong, K. Hsu, J. Huang, T. Icard, S. Jain, D. Jurafsky, P. Kalluri, S. Karamcheti, G. Keeling, F. Khani, O. Khattab, P. W. Koh, M. S. Krass, R. Krishna, R. Kuditipudi, et al., On the opportunities and risks of foundation models, CoRR abs/2108.07258 (2021). arXiv:2108.07258, doi:https://doi.org/10.48550/arXiv.2108.07258.
  • SS [29] L. Weidinger, M. Rauh, N. Marchal, A. Manzini, L. A. Hendricks, J. Mateos-Garcia, S. Bergman, J. Kay, C. Griffin, B. Bariach, et al., Sociotechnical safety evaluation of generative ai systems, arXiv preprint arXiv:2310.11986 (2023). doi:https://doi.org/10.48550/arXiv.2310.11986.
  • SS [30] D. Kang, D. Raghavan, P. Bailis, M. Zaharia, Model assertions for monitoring and improving ml models, Proceedings of Machine Learning and Systems 2 (2020) 481–496. doi:https://doi.org/10.48550/arXiv.2003.01668.
  • SS [31] Z. Yuan, Z. Xiong, Y. Zeng, N. Yu, R. Jia, D. Song, B. Li, Rigorllm: Resilient guardrails for large language models against undesired content, arXiv preprint arXiv:2403.13031 (2024). doi:https://doi.org/10.48550/arXiv.2403.13031.
  • SS [32] Z. Chu, Y. Wang, L. Li, Z. Wang, Z. Qin, K. Ren, A causal explainable guardrails for large language models, arXiv preprint arXiv:2405.04160 (2024). doi:https://doi.org/10.48550/arXiv.2405.04160
Appendix B: QAC Scores for the Selected Studies
Selected Study (SS) No. QAC1 QAC2 QAC3 QAC4 QAC5
SS-1 2 2 2 2 2
SS-2 4 5 5 5 4
SS-3 4 3 2 3 2
SS-4 3 3 3 3 4
SS-5 2 4 3 3 5
SS-6 1 5 4 3 5
SS-7 5 4 4 4 3
SS-8 5 4 4 5 3
SS-9 0 1 2 4 3
SS-10 0 2 2 3 4
SS-11 3 3 1 3 2
SS-12 3 2 3 3 2
SS-13 4 3 3 4 3
SS-14 5 4 4 5 4
SS-15 5 4 3 4 5
SS-16 5 4 4 5 4
SS-17 5 4 3 5 4
SS-18 5 4 4 5 5
SS-19 4 3 2 4 3
SS-20 5 4 4 5 4
SS-21 5 4 4 4 5
SS-22 5 5 4 5 4
SS-23 5 4 4 5 4
SS-24 5 4 2 2 3
SS-25 5 4 4 4 5
SS-26 5 4 4 5 4
SS-27 5 4 5 5 4
SS-28 5 4 4 3 4
SS-29 4 4 4 3 4
SS-30 5 4 4 4 5
SS-31 5 5 4 4 5
SS-32 5 4 4 4 5

### v4

@@ IV-A9 Customizability
The traceability attribute of guardrails tracks and records the origins, processes, and decision paths, such as input and output of FMs, external tools, etc. [19]. It involves maintaining detailed logs and records that can be audited to understand how decisions are made. For example, in a customer service chatbot, traceability ensures that every recommendation can be traced back to the data sources and algorithms used. This provides a clear audit trail for transparency and accountability. Traceability also aids in identifying the root causes of issues to enable timely and accurate troubleshooting and improvement [39], and helps in maintaining user trust and regulatory requirements [13, 7]. Additionally, comprehensive documentation of data sources and model modifications better support effective auditing and compliance checking [9].
@@ IV-B1 Actions
TABLE II: A Mapping of Agent Targets to Guardrail Actions
@@ IV-B2 Targets
Guardrail actions can be applied to various targets across multiple layers, including both pipelines and artifacts. Some guardrails are applied to the entire pipeline (including prompts, intermediate results, and final results), while others focus on specific artifacts (e.g., goals, context, reasoning, plans, memory, tools, knowledge bases, other agents, FMs). Table II summarizes agent targets and their corresponding guardrail actions. These combinations are derived from a synthesis of findings and analytical inferences based on selected studies. Moreover, they serve as illustrative examples not exhaustive or definitional, and intended to provide practical insights for potential application.
@@
  • Tools: Overseeing the proper use of tools by agents, including implementing access controls, restricting tool capabilities, and detecting potential vulnerabilities [50, 22].
@@
  • Execution Time: Managing execution time is crucial for ensuring optimal performance and timely responses for FM-based agents. Guardrails can enforce execution cut-offs for operations exceeding predefined thresholds, ensuring that agents prioritize responsiveness while avoiding excessive resource usage. For example, in conversational agents, guardrails can halt long-running processes to provide users with timely answers rather than delaying responses indefinitely.
@@ IV-B3 Rules
IV-B4 Applicability Scope
The applicability scope of guardrails in FM-based agents ranges from industry regulations and standards to individual preferences. Industry-level regulations and standards provide the broader regulatory framework within which FM-based agents must operate. Guardrails designed to comply with these regulations guarantee that the system adheres to industry best practices and legal requirements [13]. They facilitate simpler auditing and certification processes, ensuring the agent remains compliant with evolving regulatory landscapes.
@@ IV-B5 Modality
Multimodal guardrails address the combined risks of handling multiple data types. They synchronize protections across different data types, ensuring comprehensive security and compliance [26]. For example, a system that generates text based on image inputs must ensure accurate and ethical representation of the image content. This requires advanced cross-modal analysis and validation techniques to ensure the system operates reliably and ethically across all data types [28].
@@ V Reference Architecture for Designing Multi-layered Runtime Guardrails of Agents
Figure 3 illustrates the proposed reference architecture for multi-layered runtime guardrails of FM-based agents. It consists of four key components: (i) the external environment, (ii) agent components, (iii) built-in multi-layered runtime guardrails, and (iv) AgentOps infrastructure. This architecture is derived from our SLR findings (gaps and recommendations in the selected studies) and builds on empirical methodologies proposed in [66]. However, operational validation of the architecture lies beyond this paper’s scope. In the following sub-sections, we describe the components and mechanisms of this architecture and demonstrate how guardrails can be effectively implemented in FM-based agents.
@@ V-A External Environment:
The external environment refers to all entities interacting with the agent at runtime, including users, other agents, external tools, and knowledge bases. Users provide goals and contextual inputs that shape the agent’s objectives. To achieve user goals, the agent may utilize context detected in the external environment and interact with other agents, specialized tools, and extensive knowledge bases to perform complex tasks.
@@ V-B Agent Components
  • Context Engine: The context engine processes multimodal context data from the external environment to enrich the user prompt, helping FMs better understand user goals. A prompt may contain elements such as goals and context. Instead of waiting for users’ instructions, the agent can also proactively make suggestions based on the context it detects, such as screen recordings, eye tracking data, gestures, and document annotations [2].
@@
  • Memory: The memory component in this architecture stores relevant information from prior interactions, plans, and results. This accumulated knowledge supports continuous learning, enabling the agent to refine its strategies and improve capabilities and skills over time, thereby improving accuracy and minimizing errors.
@@ V-C Multi-layered Runtime Guardrails
Building on the Swiss Cheese Model, we design multi-layered runtime guardrails for FM-based agents, structured around the dimensions of quality attributes, pipelines, and artifacts specified in the taxonomy. Here, we operationalize the target mappings (discussed earlier in subsubsection IV-B2) to illustrate their implementation within the proposed reference architecture. In this architecture, each ‘cheese slice’ represents a protective layer within the agent system, addressing quality attributes, pipeline stages, and/or specific artifacts, such as one concerning privacy guardrails for prompts or security guardrails for tools. While each layer contains holes (i.e., potential gaps or weaknesses), where risks might slip through, these holes are positioned differently across layers. Gaps in one layer are often covered by another; thus, even if one layer fails, another can catch and mitigate the issue.
From the perspective of quality attributes (discussed in subsection IV-A), guardrails can be designed to ensure accuracy, efficiency, privacy, security, safety, fairness, compliance. From the pipelines perspective, guardrails can be applied at multiple stages: the user prompts, intermediate results during workflow executions, and final results generated by the agent.
@@
  • Guardrails for workflows: Handle the exceptions during the workflow executions by implementing mechanisms like force-failing a step or retrying a tool call [63]
@@
  • Guardrails for FMs: Enforce boundaries on FM’s non-deterministic outputs by applying modifications/flags
@@ VI Threats to Validity
Moreover, our definitions and categorizations may not capture all relevant aspects of guardrails in FM-based agents. To mitigate this threat, we validated the taxonomy through extensive literature review and expert feedback. However, this introduces a risk of producing biased results that address only experts’ needs, as the people involved in the feedback process have extensive experience in the AI and software engineering domains. Our review protocols helped us to reduce such bias.
We prepared a guardrails taxonomy and conducted a comparative analysis of its components to help the reader better understand their design and evaluation. We critically examined the strength and consistency of relationships in the selected studies to develop a reliable taxonomy and reference architecture for designing built-in multi-layered runtime guardrails. Finally, we draw conclusions; nonetheless, the generalizability of guardrails to different contexts and types in FM-based agents remains a potential limitation. Specific adaptations might be necessary for certain systems, such as those used in healthcare or financial organizations. Additionally, the rapid evolution of FM-based agents and their associated guardrails may lead to parts of our findings becoming outdated. To address this, continual re-evaluation and refinement of the proposed taxonomy and reference architecture will be necessary over time.
@@ VII Conclusion and Future Work
This paper presents a comprehensive taxonomy of guardrails, derived from the results of an SLR, to advance the understanding of runtime guardrails for FM-based agents. The taxonomy identifies 14 essential qualities for designing runtime guardrails in FM-based agents, addressing RQ1 in subsection IV-A. Subsequently, we outline design options for runtime guardrails, including actions, targets, employed rules, applicability scope, modality, and underlying models, addressing RQ2 in subsection IV-B. We then propose a multi-layered runtime guardrail architecture based on the taxonomy, SLR findings, and the empirically grounded reference architecture approach [66]. This novel guardrail architecture better addresses the unique challenges associated with FM-based agents, as discussed in section V, thus directly answering RQ3. In future work, we plan to develop guardrail services for a scientific agent platform. These services will implement the proposed reference architecture and integrate the various design options outlined in the taxonomy.